πΎπ‘Ž (𝑁𝐻4 +) = 10βˆ’9 = 𝑁𝐻3 [𝐻3𝑂 +] [𝑁𝐻4 +] = π‘₯ βˆ— π‘₯ 0.1 βˆ’ π‘₯ β‰ˆ π‘₯ 2 0.1 β‡’ π‘₯ =


1
Expert's answer
2021-06-25T05:43:46-0400

Ans:-

"\to 𝐾𝑎 (𝑁𝐻4 ^+)={NH_3}{[H_3O^+][NH_4^+]}" for weak electrolyte


"10^{-9}=x^2\times{(0.1-x)}"


We know that "\ \ \ \ 0.1 − 𝑥 ≈ 0.1"

"10^{-9}=x^2\times{0.1}\Rightarrow x=10^{-4}"